OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 8 inches. What is the volume of water in the vase? 26π in3 16π in3 24π in3 32π in3

OpenStudy (anonymous):

@Shane_B

OpenStudy (anonymous):

@Mathmuse

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

i saw that one but i still got the wrong answer

OpenStudy (anonymous):

oh ok

OpenStudy (shane_b):

Then it would probably help if you showed your work so we can find the mistake :)

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

ok, so this is what i did: I took tha equation pir^2h, and plugged in the numbers. So it loked liked this: (3.14)*(4)(8). Then i took the equation 4/3pir^3, and plugged in the numbers: 4/3*(3.14)(8). Then multiplied that by 6 since there were 6 marbles. Then i subtracted them: 100.48*-200.94=-100.46. See.. totally wrong

OpenStudy (shane_b):

Why is there an 8 here? 4/3*(3.14)(8) The radius of the sphere is just 1.

OpenStudy (anonymous):

but the heiight is 8 iches

OpenStudy (anonymous):

i guess i dont get it lol

OpenStudy (shane_b):

But you're mixing the sphere equation with the cylinder equation.

OpenStudy (anonymous):

i have to frigure out what the sheres volumes are so i can subtract that from the totalyvolume

OpenStudy (anonymous):

*total volume

OpenStudy (anonymous):

that is what i tried to do and i got it wayy wrong

OpenStudy (anonymous):

why

OpenStudy (anonymous):

click 24

OpenStudy (anonymous):

k

OpenStudy (anonymous):

maybe thats the answer

OpenStudy (anonymous):

maybe lol

OpenStudy (anonymous):

want churritos

OpenStudy (anonymous):

YESS! i love those things!

OpenStudy (anonymous):

<------me 2 haha

OpenStudy (anonymous):

haha aww! bring me some! :P

OpenStudy (anonymous):

here --------_______------ hahahaha

OpenStudy (anonymous):

hha YAY!

OpenStudy (shane_b):

The volume of the cylinder will be:\[V{cyl}=\pi r^2h=\pi (2)^2(8)=32\pi \space in^2\]\[V_{marble}=\frac{4}{3}\pi r^3=\frac{4}{3}\pi(1)^3=\frac{4}{3}\pi\]Now multiply the volume of a marble times 6 to get the volume of all of them:\[(6)\frac{4}{3} \pi=8\pi \space in^2\]\[32\pi - 8\pi = 24\pi\]

OpenStudy (anonymous):

seee i was right

OpenStudy (shane_b):

Well, you had a 25% chance of being right at a guess :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks to both of you guys.. and i kind of get it now! yay!

OpenStudy (shane_b):

Great, you're welcome :)

OpenStudy (anonymous):

bye guys net ?

OpenStudy (anonymous):

i mean next